# Minimal pole representation for spectral functions

### v1

@@ III.3.2 Sub-Ohmic bath
In this section, we examine a sub-Ohmic bath ($s=0.5$) at $T=50\;{\rm K}$. The spectral function, which is shown in Fig. 6(a), diverges at $\omega=0$, but its integral remains finite, ensuring the existence of the BCF. However, the real part of $C(t)$, which is shown in Fig. 6(b), exhibits an extremely slowly decaying tail.
@@
Figure 6: Comparison for a sub-Ohmic bath at $T=50\,{\rm K}$. Eight terms are used for all approximations. (a) Exact spectral function, which has a divergence at $\omega=0$. (b) Exact bath correlation function, which exhibits an extremely slowly decaying tail for the real part. The insets show the performance of different methods in a zoomed-in region. Mean error for the approximated interval and the long tail are shown in subfigures (c) and (d), respectively.
@@ III.4 Decomposing structured spectra for real-time evolution
Finally, we evaluate the performance of the methods on a structured spectrum derived from the spin-boson model, a two-level system coupled to a bosonic bath.
@@
The input data consists of the spectral density $J(\omega)$ sampled on a dense uniform grid. As a preprocessing step, the AAA algorithm is applied to the effective spectrum at $T=300\;{\rm K}$, prioritizing accuracy over the number of poles. The error tolerance is set to $10^{-8}$, resulting in 704 poles $\{A_{l}^{({\rm AAA})},\xi_{l}^{({\rm AAA})}\}$, evenly distributed between the upper and lower half-planes. These poles are treated as the exact solution and serve as the reference for all methods compared in this section. In this case, the contour integral in Eq. (10) for MPM can be evaluated analytically using $h_{k}=\sum_{l}A_{l}^{{}^{\prime}({\rm AAA})}(\xi_{l}^{{}^{\prime}({\rm AAA})})^{k}$ [Eq. (11)] instead of a numerical integral, thereby improving computational efficiency.
@@
We present the results in Fig. 7, where $M$ is fixed at 100. For AAA, the number of poles to be recovered is set to $2M$, as it captures poles across the entire complex plane. For ESPRIT, the parameters $t_{c}=20000\;({\rm eV})^{-1}$ and $N_{t}=5000$ are used. Issues arise for both the AAA and ESPRIT algorithms. On one hand, the original AAA method consistently recovers some poles on the real axis, leading to non-causal jumps in the spectrum. Denoting these poles as $\{A_{l}^{(\rm real)},\xi_{l}^{(\rm real)}\}$ and their count as $M_{\rm real}$, these spurious poles introduce a constant term in $C(t)$, degrading accuracy in both the real-frequency and real-time domains. To ensure a reliable real-time evolution, the method should be modified to remove these poles. However, as shown in Fig. 7, even after correction, both the approximated spectrum and the BCF exhibit some deviations from the exact results. On the other hand, ESPRIT becomes unstable for some values of the parameter $L$. Although this issue can be mitigated by carefully tuning $L$, allowing ESPRIT to approximate the BCF similarly to MPM, we note that it consistently recovers some poles in the upper half plane. This leads to a divergent amplitude of the BCF at long times, even though the exact BCF exhibits decay. To eliminate the asymptotic instability, these non-physical poles must be removed [66]. In practice, this is done by first obtaining $M$ poles from ESPRIT, discarding the unphysical ones, and then performing a least-squares fit to determine the weights of the remaining poles. Unsurprisingly, we find that once some poles are disregarded, ESPRIT immediately loses error control, as evidenced by its deviation from the exact results in Figs. 7(b) and (c).
@@
In contrast, MPM does not suffer from these issues. Its poles are automatically placed in the lower half-plane, ensuring that the approximation remains stable and well-controlled. As a result, MPM provides the most accurate approximation in both the time and frequency domains. Furthermore, in this case, MPM (M1) and MPM (M2) yield nearly identical results due to the rapid decay of the long tails in $A(\omega)$. Consequently, we present only MPM (M1) results throughout this section, with $\omega_{p}$ fixed at 0.04 eV and $k_{\rm max}$ set to 3000.
@@ IV Conclusions
Our paper is accompanied by an open-source software implementation of the proposed method, written in the programming language PYTHON [67, 68].
@@ Acknowledgements.
We thank Michael Thoss, Samuel Rudge, Salvatore Gatto, and James P. F. LeBlanc for helpful discussions. We are especially grateful to Michael Thoss, Samuel Rudge, Salvatore Gatto, Eran Rabani and Bokang Hou for providing the data used for the structured spectra in Sec. III.4. This material is based upon work supported by the National Science Foundation under Grant No. 2310182. Real-time aspects have been supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award No. DE-SC0022088.
@@ Appendix A Prony-like Problem and ESPRIT Algorithm
Many numerical methods have been proposed to estimate $M$, $R_{l}$, and $z_{l}$ in Eq. (46). These include the Prony approximation method [69, 70], the Matrix Pencil Method [31, 71], and the Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) [32, 62]. Due to its superior performance [53], we focus exclusively on the ESPRIT algorithm in this paper.
@@
where $U$ and $W$ are unitary matrices of dimensions $(N-L)\times(N-L)$ and $(L+1)\times(L+1)$, respectively. The matrix $\Sigma$ is a rectangular diagonal matrix of size $(N-L)\times(L+1)$, where diagonal elements are ordered as $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{L+1}\geq 0$. In practice, $L$ is typically chosen between $N/3$ and $N/2$ to minimize variance [71]. Throughout our implementation, we set $L=2N/5$ unless otherwise stated.
@@
In cases where the sampled function is matrix-valued, a generalization of ESPRIT to matrix-valued functions [22] can be applied.

### v2

@@ III.3.2 Sub-Ohmic bath
In this section, we examine a sub-Ohmic bath ($s=0.5$) at $T=50\;{\rm K}$ ($\beta\approx 153\;{\rm fs}$). The spectral function, which is shown in Fig. 6(a), diverges at $\omega=0$, but its integral remains finite, ensuring the existence of the BCF. However, the real part of $C(t)$, which is shown in Fig. 6(b), exhibits an extremely slowly decaying tail.
@@
Figure 6: Comparison for a sub-Ohmic bath at $T=50\,{\rm K}$ ($\beta\omega_{c}\approx 1.44$). Eight terms are used for all approximations. (a) Exact spectral function, which has a divergence at $\omega=0$. (b) Exact bath correlation function, which exhibits an extremely slowly decaying tail for the real part. The insets show the performance of different methods in a zoomed-in region. Mean error for the approximated interval and the long tail are shown in subfigures (c) and (d), respectively.
@@ III.4 Decomposing structured spectra for real-time evolution
Finally, we evaluate the performance of the methods on a structured spectrum derived from exciton–phonon couplings in quantum dots Jasrasaria and Rabani (2023, 2021), which plays a key role in understanding nonradiative exciton dynamics in these systems.
@@
The input data consists of the spectral density $J(\omega)$ sampled on a dense uniform grid. To enable flexible sampling at arbitrary frequency points, the AAA algorithm is applied as a preprocessing step to the effective spectrum at $T=300\;{\rm K}$, with accuracy prioritized over the number of poles. The error tolerance is set to $10^{-8}$, resulting in 704 poles $\{A_{l}^{({\rm AAA})},\xi_{l}^{({\rm AAA})}\}$, evenly distributed between the upper and lower half-planes. These poles are treated as the exact solution and serve as the reference for all methods compared in this section. In this case, the contour integral in Eq. (10) for MPM can be evaluated analytically using $h_{k}=\sum_{l}A_{l}^{{}^{\prime}({\rm AAA})}(\xi_{l}^{{}^{\prime}({\rm AAA})})^{k}$ [Eq. (11)] instead of a numerical integral, thereby improving computational efficiency.
@@
We present the results in Fig. 7, where $M$ is fixed at 100. For AAA, the number of poles to be recovered is set to $2M$, as it captures poles across the entire complex plane. For ESPRIT, the parameters $t_{c}=20000\;{\rm eV}^{-1}\approx 8.27\times 10^{4}\;{\rm fs}$ and $N_{t}=5000$ are used. Issues arise for both the AAA and ESPRIT algorithms. On one hand, the original AAA method consistently recovers some poles on the real axis, leading to non-causal jumps in the spectrum. Denoting these poles as $\{A_{l}^{(\rm real)},\xi_{l}^{(\rm real)}\}$ and their count as $M_{\rm real}$, these spurious poles introduce a constant term in $C(t)$, degrading accuracy in both the real-frequency and real-time domains. To ensure a reliable real-time evolution, the method should be modified to remove these poles. However, as shown in Fig. 7, even after correction, both the approximated spectrum and the BCF exhibit some deviations from the exact results. On the other hand, ESPRIT becomes unstable for some values of the parameter $L$. Although this issue can be mitigated by carefully tuning $L$, allowing ESPRIT to approximate the BCF similarly to MPM, we note that it consistently recovers some poles in the upper half plane. This leads to a divergent amplitude of the BCF at long times, even though the exact BCF exhibits decay. To eliminate the asymptotic instability, these non-physical poles must be removed Dunn, Tempelaar, and Reichman (2019). In practice, this is done by first obtaining $M$ poles from ESPRIT, discarding the unphysical ones, and then performing a least-squares fit to determine the weights of the remaining poles. Unsurprisingly, we find that once some poles are disregarded, ESPRIT immediately loses error control, as evidenced by its deviation from the exact results in Figs. 7(b) and (c).
@@
In contrast, MPM does not suffer from these issues. Its poles are automatically placed in the lower half-plane, ensuring that the approximation remains stable and well-controlled. As a result, MPM provides the most accurate approximation in both the time and frequency domains. Furthermore, in this case, MPM (M1) and MPM (M2) yield nearly identical results due to the rapid decay of the long tails in $A(\omega)$. Consequently, we present only MPM (M1) results throughout this section, with $\omega_{p}$ fixed at $0.04\;{\rm eV}\approx 9.67\times 10^{-3}\;{\rm fs}^{-1}$ and $k_{\rm max}$ set to 3000.
@@ IV Conclusions
Our paper is accompanied by an open-source software implementation of the proposed method, written in the programming language PYTHON Zhang et al. (2025a, b).
@@ Acknowledgements.
We thank Michael Thoss, Samuel Rudge, Salvatore Gatto, Eran Rabani, Bokang Hou and James P. F. LeBlanc for helpful discussions. This material is based upon work supported by the National Science Foundation under Grant No. 2310182. Real-time aspects have been supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award No. DE-SC0022088.
DATA AVAILABILITY
The data and the code that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Prony-like Problem and ESPRIT Algorithm
Many numerical methods have been proposed to estimate $M$, $R_{l}$, and $z_{l}$ in Eq. (47). These include the Prony approximation method Beylkin and Monzón (2005, 2010), the Matrix Pencil Method Hua and Sarkar (1990); Sarkar and Pereira (1995), and the Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) Roy and Kailath (1989); Potts and Tasche (2013). Due to its superior performance Takahashi et al. (2024), we focus exclusively on the ESPRIT algorithm in this paper.
@@
where $U$ and $W$ are unitary matrices of dimensions $(N-L)\times(N-L)$ and $(L+1)\times(L+1)$, respectively. The matrix $\Sigma$ is a rectangular diagonal matrix of size $(N-L)\times(L+1)$, where diagonal elements are ordered as $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{L+1}\geq 0$. In practice, $L$ is typically chosen between $N/3$ and $N/2$ to minimize variance Sarkar and Pereira (1995). Throughout our implementation, we set $L=2N/5$ unless otherwise stated.
@@
In cases where the sampled function is matrix-valued, a generalization of ESPRIT to matrix-valued functions Zhang, Yu, and Gull (2024) can be applied.